A Sound note is produced by a ringing bell and the echo of the note from a nearby wall is received 0.5secs later. If the frequency of the note is 400hz and its wavelength is 1m. Calculate the distance between the bell and the wall
1
Expert's answer
2020-04-21T11:00:08-0400

The speed of wave:


"v=f\lambda=(400)(1)=400\frac{m}{s}"

The distance between the bell and the wall


"d=0.5vT=0.5(400)(0.5)=100\ m"